# A binary neutron star merger search pipeline powered by deep learning

### v1

@@ Abstract
Gravitational waves are now routinely detected from compact binary mergers, with binary neutron star mergers being of note for multi-messenger astronomy as they have been observed to produce electromagnetic counterparts. Novel search pipelines for these mergers could increase the combined search sensitivity, and could improve the ability to detect real gravitational wave signals in the presence of glitches and non-stationary detector noise. Deep learning has found success in other areas of gravitational wave data analysis, but a sensitive deep learning-based search for binary neutron star mergers has proven elusive due to their long signal length. In this work, we present a deep learning pipeline for detecting binary neutron star mergers. By training a convolutional neural network to detect binary neutron star mergers in the signal-to-noise ratio time series, we concentrate signal power into a shorter and more consistent timescale than strain-based methods, while also being able to train our network to be robust against glitches. We compare our pipeline’s sensitivity to the three offline detection pipelines using injections in real gravitational wave data, and find that our pipeline has a comparable sensitivity to the current pipelines below the 1 per 2 months detection threshold. Furthermore, we find that our pipeline can increase the total number of binary neutron star detections by 12% at a false alarm rate of 1 per 2 months. The pipeline is also able to successfully detect the two binary neutron star mergers detected so far by the LIGO-Virgo-KAGRA collaboration, GW170817 and GW190425, despite the loud glitch present in GW170817.
@@ I Introduction
Gravitational waves (GW) from compact binary coalescences (CBCs) are now regularly detected by ground-based laser interferometers, with the LIGO [1] and Virgo [2] interferometers detecting over 90 CBCs in the first three observing runs [3, 4, 5]. Most of these CBCs are binary black hole (BBH) mergers, with only two confirmed binary neutron star (BNS) mergers [6, 7], and two confirmed neutron star - black hole (NSBH) mergers by the end of the third observing run (O3) of the LIGO-Virgo-KAGRA collaboration. The first detected BNS merger, GW170817, is notable for its heralding of a new age of multi-messenger astronomy with gravitational waves as a messenger [6, 8]. A gamma-ray burst was serendipitously detected from the merger [9], as well as a kilonova and X-ray counterpart from follow-up observations [8, 10]. These observations provided a unique measurement of the Hubble constant [11], and constraints on the neutron star equation of state [12, 13]. Further observations of BNS mergers could further constrain the Hubble constant and resolve the Hubble tension, and potentially reveal a link between BNS mergers and other transient signals, such as fast radio bursts. As the interferometers improve in sensitivity, and new interferometers such as KAGRA [14] come online, the possibility of making more multi-messenger detections warrants the development of new CBC pipelines to search for them.
@@
Despite the success of the current pipelines at detecting CBCs, it is worthwhile investigating new detection methods for several reasons. Firstly, the overall search for CBCs benefits from multiple search pipelines using unique search methods [21]. Unique search methods provide the possibility for the detection of events that would have been missed by other search methods, and joint detections with other searches provide supporting evidence that an event is a genuine CBC. Secondly, an area of ongoing research is the mitigation of non-Gaussian transient noise artefacts (glitches), which can produce high-SNR triggers that pipelines should avoid producing alerts on. A key challenge is the exclusion of glitches from analysis, without also excluding actual CBC signals from producing alerts [6]. As the detection rate of CBCs and the rate of instrumental glitches have both increased over time [4, 5], CBC events contaminated by glitches will likely become more frequent in the future. A detection method that can successfully identify signals while minimising the effect of glitches, as well as correctly identifying signals contaminated with glitches would be ideal. With these stipulations, a deep learning-based detection method is a logical choice for investigation [22].
@@
Deep learning has found success as a useful tool for improving the accuracy or latency of several areas of gravitational wave data analysis (see e.g. [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 22]). Gravitational wave strain-based BBH detection with deep learning has been shown to be effective, and potentially able to reach the sensitivity of the matched filtering detection pipelines in real detector noise [30, 31, 32]. However, compared to BBH detection, there are additional challenges introduced when applying deep learning to lower mass signals like BNS mergers. Strain-based BNS detection methods face the issue that at current sensitivity, BNS mergers are present for $\mathcal{O}(100~{}\text{s})$ in detector data, meaning the signal power that accumulates at a detector is significantly more spread out compared to a BBH merger with equivalent SNR. A strain-based BNS detection method [33, 34, 35, 36, 37] has to either lose signal power by truncating the input window, or through other approximations made in pre-processing, which limits the achievable sensitivity. Spectrogram-based detection methods [38, 39] face the same issue. Consequently, a deep learning approach for BNS detection that can match the sensitivity of the matched filtering detection pipelines has yet to be demonstrated.
@@
In this work, we investigate the use of a neural network (NN) based search pipeline for the detection of BNS mergers in the SNR time series produced by matched filtering. The advantage of detecting in the SNR time series is that CBC signal power is condensed relative to in the strain, which is especially beneficial for the longer-duration BNS mergers. SNR time series are also a readily available data product from the matched filtering pipelines, making online implementation relatively straightforward. This work is further motivated by [22], where we found that BBH detection with SNR time series produced promising sensitivity results, especially towards lower BBH masses. We train our NN on LIGO Hanford (H1) and LIGO Livingston (L1) real detector noise from the third observing run to ensure it is robust against real glitches. By characterising our search pipeline’s ranking statistic on past data and performing searches on the O3 injection set from the third gravitational wave transient catalogue (GWTC-3) [5], we show that our NN can match the sensitivity of the current detection pipelines and can detect the two real BNS events, GW170817 and GW190425.
@@
The structure of the remainder of this work is as follows. In Sec. II we cover how we implement matched filtering, select our detector noise, and how we generate our template bank and training datasets. In Sec. III we cover the high-level architecture of our neural network and its training and validation. The method we use to run our search pipeline and assign false alarm rates is presented in Sec. IV. In Sec. V we present the performance of our search pipeline using an injection set from the GWTC-3 offline analyses, as well as the pipeline’s detection of the two real BNS events. In Sec. VI we summarise the findings of this work, discuss their implications, and discuss potential future improvements.
@@ II.1 Matched filtering
Matched filtering is a signal processing technique commonly employed in gravitational wave research, as it is the optimal detection method for modelled signals in stationary Gaussian noise [40]. It is the process of cross-correlating a signal template $s$ with incoming detector data $h$, and produces a signal-to-noise ratio (SNR) time series $\rho(t)$ [40, 41]:
$\rho^{2}(t)=\frac{z(t)}{\langle s|s\rangle}\,,$ (1)
where $\langle s|s\rangle$ is the noise-weighted inner product of the template and $z(t)$ is the matched filter
@@
where $S_{n}(f)$ is the estimated one-sided power spectral density (PSD) of the detector noise and a tilde represents the Fourier transform of the template or data.
@@ II.3 Noise data selection and glitch identification
To generate our datasets, we fetched O3 public data in 1 week chunks [48], and only fetched segments of data where both LIGO Hanford and LIGO Livingston are online, as these were the most sensitive detectors during O3. While a single-detector search pipeline would be useful for when one of the LIGO detectors is down, we currently only consider the two-detector case as this configuration is more sensitive, and handling arbitrary pairs of interferometers adds complexity. We choose to process data in segments of 1,024 seconds, so segments shorter than this duration are excluded. We also exclude segments of data containing GW signals from our training datasets, as identified by the GWTC catalogues [3, 4, 5]. Neither of these restrictions significantly reduces the amount of data available. For the first week of O3, these criteria lead to a duty factor of $\sim 55\%$, which is comparable to the combined duty factor of the two LIGO detectors during O3a. Based on the selected segments of data, we then produce a list of valid integer GPS times with the only requirements being that the merger time is at least 100 seconds from the start of the segment (slightly longer than the longest possible waveform), and at least 24 seconds from the end of the segment. These valid GPS times are used for assigning training and validation samples a merger time. The only pre-processing we apply to the data is downsampling it to 2,048 Hz. We compute the PSD for each week of data, which is used for normalising the SNR time series. We use the Welch method to compute the PSD with a 4-second window for each 1,024-second segment of data. These PSDs are then averaged together for a mean PSD for each week of data.
@@ II.4 Training dataset construction
For our training dataset, we next assign a random GPS time from the first week of O3 to each event. To ensure we have enough unique noise realisations, we use a different random GPS time for each interferometer. We also specify the probability for a sample to contain a glitch. Based on the number of glitches found by Omicron, we chose the Hanford glitch fraction to be 0.1, and the Livingston glitch fraction to be 0.15, as glitches occurred more frequently in Livingston [5]. These glitch fractions are around an order of magnitude higher than the actual rate of glitches during O3, but we specified these fractions to ensure the neural network is sufficiently trained on glitches.
@@
Strain samples with injections are created by first generating the injection waveform with the SpinTaylorT4 approximant [45]. The waveforms are then projected onto the detectors and placed with the merger at the centre of a 1024s-long noise segment based on the Hanford GPS time, with the Livingston detector signal offset based on the extrinsic parameters. The strain is then filtered with the 10 selected templates, and the resulting SNR time series are then sliced down to a 2-second time series centred on the merger. As our neural network’s input window is 1 second long, this 2-second slice of SNR time series allows the merger time to be placed randomly within the input window. However, to ensure the entire peak from the merger in both interferometers is completely contained within an SNR time series, we place the merger at least 1/8th of a second from the edges of the window.
@@
Polarisation Uniform [0, $\pi$]
Network SNR - $\geq$ 6
@@ III.1 Model architecture
Figure 1: a) The high-level architecture of our neural network. Each block represents a layer or group of layers, with the height of the block representing the number of convolution kernels and the width representing output size of the layer. The Hanford input branch is shown, and the Livingston input branch has the same architecture. b) The internal structure of each residual block. c) A legend of layers and their representation in the diagram. Residual blocks with a purple stripe contain a maximum pooling layer. During training, there is a dropout layer between the dense layers, and a sigmoid activation layer after the last dense layer.
The network is constructed such that it can be split into three sub-networks around the addition layer between the two branches. The three resulting networks (hereafter called the H, L and combiner sub-networks) can then be operated independently. Since the combiner sub-network is a simple dense network, it can be run much faster on a CPU than the H and L sub-networks that require a GPU to run efficiently. The advantage of this network architecture is that the outputs of the H and L sub-networks can be independently paired up before being input into the combiner sub-network, which aids in constructing our background to characterise our ranking statistic, as further explained in Sec. IV. One downside of this architecture is that any coincidence information between the two detectors is lost. To mitigate this, an additional input $\Delta t$ is passed to the combiner sub-network as well as the H and L sub-network outputs:
@@
During training, there are three additional features that are removed during inference. Firstly, we use a dropout layer between each dense layer in the combiner sub-network to reduce overfitting on the training set. Secondly, we use a sigmoid activation layer after the final dense layer to constrain the prediction to between 0 and 1. Removing this activation during inference has been shown to be effective at mitigating the resolution limitations of 32-bit precision [53] and allows our ranking statistic to be unbounded. Thirdly, we add a custom layer before the sigmoid activation layer which divides the output of the last dense layer by a factor of 4, which helps prevent the sigmoid layer from rounding predictions to 0 or 1 during training.
@@ III.2 Training
The network was trained using the training and validation datasets described in Sec. II.4. The training and validation samples were weighted in training if their network SNR was greater than 10 by a factor of SNR/10, as we found that the network tended to assign very low prediction values to very high SNR signals, likely because it was mistaking them for glitches. This sample weighting factor also applied to noise-only samples, and so the neural network was additionally penalised for labelling SNR time series with high SNR glitches as signals.
The network was trained with the binary crossentropy loss function using the ADAM optimiser [54], and an initial learning rate of $10^{-4}$. We use the Keras callbacks ReduceLROnPlateau and EarlyStopping to fine-tune the network further: if the validation loss did not improve in 15 epochs, the learning rate was halved, and training stopped if the network’s validation loss did not improve after 25 epochs. We created a custom metric, which we call LogAUC, for EarlyStopping to determine when the network should stop being trained. LogAUC is based on Keras’s AUC metric, which calculates the true alarm probability and false alarm probability at different thresholds (i.e. a receiver operating characteristic, or ROC curve), then computes the area under the resulting curve. Maximising the area under this curve ensures a trained model is sensitive at the tested false alarm thresholds. LogAUC simply spaces the false alarm rate thresholds logarithmically and calculates the area of the curve in log space, rather than linearly, which ensures that the sensitivity at low false alarm rates is given suitable weight during training. This is important as GW detection requires a focus on low false alarm rates assigned over many orders of magnitude. Training the network typically takes 3 hours on an NVIDIA A100 GPU.
@@ IV Search Method and False Alarm Rate Assignment
Here we describe how we run our search pipeline and compute our ranking statistic, which is used for assigning false alarm rates to events. Since we detect in the SNR time series rather than the strain, our ranking statistic involves several data reduction steps to reduce the computational requirements of the neural network. The ranking statistic for a second of data is computed with the following workflow. First, the SNR time series are computed using the entire template bank. These SNR time series are computed in clusters of 30 due to memory limitations, and are ordered by chirp mass (i.e. all templates in the cluster have a similar chirp mass). For each second of SNR time series, we use a coincident peak-finding algorithm to find the highest network SNR of the time series. A trigger is produced if a peak is found with SNR $>4$ in at least one of the detectors. Only one trigger is saved per SNR time series per second.
@@
The SNR time series with the highest network SNR trigger from each cluster is then used as the input to the neural network. The neural network then makes 16 predictions per second on the SNR time series, i.e. an inference rate of 16 Hz. The moving average of each series of predictions is then used as the ranking statistic on that second of data with that template. Using an inference rate greater than 1 Hz aids with separating glitches and high SNR noise from signal candidates, and this benefit has been noted in previous works [22, 32, 55]. A noise peak is unlikely to produce multiple high-valued predictions in a row, while a peak from a signal would consistently produce high predictions while in the input window. More information on the effect of the inference rate on sensitivity can be found in Sec. V.4. The moving average predictions from each of the template clusters are then compared, and the highest moving average prediction becomes the ranking statistic for that second. In addition to the ranking statistic, the network SNR of the trigger, the time of the peak SNR and the ID of the template that produced the trigger are saved for each second of data.
@@
To characterise our ranking statistic background, we collect the ranking statistic on one week of noise, for which we use the third week of O3. To extend our background, we perform 100 time shifts for each second of data. Time shifts are a simple method for increasing the size of a background, thus making it a more accurate distribution, without requiring an excessive amount of real detector data. Time shifts typically involve shifting the data of two interferometers by a time greater than their light travel time and then computing the ranking statistic of this new pair of data. In our case, however, time shifting the SNR time series would be computationally infeasible, as this would result in a hundredfold increase in the number of predictions our neural network would have to make. Our solution to this is to divide the network into the H, L and combiner sub-networks (as described in Sec. III), and time shift the outputs of the H and L sub-networks when collecting our background. The benefit of splitting the model this way is that only the combiner sub-network has to process 100 time shifts, and the H and L sub-networks only process the single week of background data. Since the combiner sub-network is much less computationally expensive than the H and L sub-networks, it can predict on 100 time shifted samples when running on a CPU at the same pace as the H and L models running on a GPU. This procedure is similar to that described in [56]. In our implementation, we keep the Hanford SNR time series fixed and shift the Livingston SNR time series by one second for each time shift.
@@
Figure 2: Ranking statistic distribution from 1.1 years of background data, and the corresponding false alarm rates. The orange dashed line shows the Gaussian fit used to assign false alarm rates to ranking statistics higher than the background.

### v2

@@ Abstract
Gravitational waves are now routinely detected from compact binary mergers, with binary neutron star mergers being of note for multimessenger astronomy as they have been observed to produce electromagnetic counterparts. Novel search pipelines for these mergers could increase the combined search sensitivity, and could improve the ability to detect real gravitational wave signals in the presence of glitches and nonstationary detector noise. Deep learning has found success in other areas of gravitational wave data analysis, but a sensitive deep learning-based search for binary neutron star mergers has proven elusive due to their long signal length. In this work, we present a deep learning pipeline for detecting binary neutron star mergers. By training a convolutional neural network to detect binary neutron star mergers in the signal-to-noise ratio time series, we concentrate signal power into a shorter and more consistent timescale than strain-based methods, while also being able to train our network to be robust against glitches. We compare our pipeline’s sensitivity to the three offline detection pipelines using injections in real gravitational wave data, and find that our pipeline has a comparable sensitivity to the current pipelines below the 1 per 2 months detection threshold. Furthermore, we find that our pipeline can increase the total number of binary neutron star detections by 12% at a false alarm rate of 1 per 2 months. The pipeline is also able to successfully detect the two binary neutron star mergers detected so far by the LIGO-Virgo-KAGRA collaboration, GW170817 and GW190425, despite the loud glitch present in GW170817.
@@ I Introduction
Gravitational waves (GW) from compact binary coalescences (CBCs) are now regularly detected by ground-based laser interferometers, with the LIGO [1] and Virgo [2] interferometers detecting over 90 CBCs in the first three observing runs [3, 4, 5]. Most of these CBCs are binary black hole (BBH) mergers, with only two confirmed binary neutron star (BNS) mergers [6, 7], and two confirmed neutron star - black hole (NSBH) mergers by the end of the third observing run (O3) of the LIGO-Virgo-KAGRA collaboration. The first detected BNS merger, GW170817, is notable for its heralding of a new age of multimessenger astronomy with gravitational waves as a messenger [6, 8]. A gamma-ray burst was serendipitously detected from the merger [9], as well as a kilonova and x-ray counterpart from follow-up observations [8, 10]. These observations provided a unique measurement of the Hubble constant [11], and constraints on the neutron star equation of state [12, 13]. Further observations of BNS mergers could further constrain the Hubble constant and resolve the Hubble tension, and potentially reveal a link between BNS mergers and other transient signals, such as fast radio bursts. As the interferometers improve in sensitivity, and new interferometers such as KAGRA [14] come online, the possibility of making more multimessenger detections warrants the development of new CBC pipelines to search for them.
@@
Despite the success of the current pipelines at detecting CBCs, it is worthwhile investigating new detection methods for several reasons. First, the overall search for CBCs benefits from multiple search pipelines using unique search methods [21]. Unique search methods provide the possibility for the detection of events that would have been missed by other search methods, and joint detections with other searches provide supporting evidence that an event is a genuine CBC. Second, an area of ongoing research is the mitigation of non-Gaussian transient noise artefacts (glitches), which can produce high-SNR triggers that pipelines should avoid producing alerts on. A key challenge is the exclusion of glitches from analysis, without also excluding actual CBC signals from producing alerts [6]. As the detection rate of CBCs and the rate of instrumental glitches have both increased over time [4, 5], CBC events contaminated by glitches will likely become more frequent in the future. A detection method that can successfully identify signals while minimizing the effect of glitches, as well as correctly identifying signals contaminated with glitches would be ideal. With these stipulations, a deep learning-based detection method is a logical choice for investigation [22].
@@
Deep learning has found success as a useful tool for improving the accuracy or latency of several areas of gravitational wave data analysis (see e.g. [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 22]). Gravitational wave strain-based BBH detection with deep learning has been shown to be effective, and potentially able to reach the sensitivity of the matched filtering detection pipelines in real detector noise [30, 31, 32]. However, compared to BBH detection, there are additional challenges introduced when applying deep learning to lower mass signals like BNS mergers. Strain-based BNS detection methods face the issue that at current sensitivity, BNS mergers are present for $\mathcal{O}(100~{}\text{s})$ in detector data, meaning the signal power that accumulates at a detector is significantly more spread out compared to a BBH merger with equivalent SNR. A strain-based BNS detection method [33, 34, 35, 36, 37] has to either lose signal power by truncating the input window, or through other approximations made in preprocessing, which limits the achievable sensitivity. Spectrogram-based detection methods [38, 39] face the same issue. Consequently, a deep learning approach for BNS detection that can match the sensitivity of the matched filtering detection pipelines has yet to be demonstrated.
@@
In this work, we investigate the use of a neural network (NN) based search pipeline for the detection of BNS mergers in the SNR time series produced by matched filtering. The advantage of detecting in the SNR time series is that CBC signal power is condensed relative to in the strain, which is especially beneficial for the longer-duration BNS mergers. SNR time series are also a readily available data product from the matched filtering pipelines, making online implementation relatively straightforward. This work is further motivated by [22], where we found that BBH detection with SNR time series produced promising sensitivity results, especially toward lower BBH masses. We train our NN on LIGO Hanford (H1) and LIGO Livingston (L1) real detector noise from the third observing run to ensure it is robust against real glitches. By characterizing our search pipeline’s ranking statistic on past data and performing searches on the O3 injection set from the third gravitational wave transient catalog (GWTC-3) [5], we show that our NN can match the sensitivity of the current detection pipelines and can detect the two real BNS events, GW170817 and GW190425.
@@
The structure of the remainder of this work is as follows. In Sec. II we cover how we implement matched filtering, select our detector noise, and how we generate our template bank and training datasets. In Sec. III we cover the high-level architecture of our neural network and its training and validation. The method we use to run our search pipeline and assign false alarm rates is presented in Sec. IV. In Sec. V we present the performance of our search pipeline using an injection set from the GWTC-3 offline analyses, as well as the pipeline’s detection of the two real BNS events. In Sec. VI we summarize the findings of this work, discuss their implications, and discuss potential future improvements.
@@ II.1 Matched filtering
Matched filtering is a signal processing technique commonly employed in gravitational wave research, as it is the optimal detection method for modeled signals in stationary Gaussian noise [40]. It is the process of cross-correlating a signal template $s$ with incoming detector data $h$, and produces a signal-to-noise ratio (SNR) time series $\rho(t)$ [40, 41]:
$\rho(t)=\frac{|z(t)|}{\sqrt{\langle s|s\rangle}}\,,$ (1)
where $\langle s|s\rangle$ is the noise-weighted inner product of the template and $z(t)$ is the complex matched filter
@@
where $S_{n}(f)$ is the estimated one-sided power spectral density (PSD) of the detector noise, $f_{low}$ and $f_{high}$ are the low and high frequency cutoffs used, and a tilde represents the Fourier transform of the template or data.
@@ II.3 Noise data selection and glitch identification
To generate our datasets, we fetched O3 public data in 1 week chunks [48], and only fetched segments of data where both LIGO Hanford and LIGO Livingston are online, as these were the most sensitive detectors during O3. While a single-detector search pipeline would be useful for when one of the LIGO detectors is down, we currently only consider the two-detector case as this configuration is more sensitive, and handling arbitrary pairs of interferometers adds complexity. We choose to process data in segments of 1,024 seconds, so segments shorter than this duration are excluded. We also exclude segments of data containing GW signals from our training datasets, as identified by the GWTC catalogs [3, 4, 5]. Neither of these restrictions significantly reduces the amount of data available. For the first week of O3, these criteria lead to a duty factor of $\sim 55\%$, which is comparable to the combined duty factor of the two LIGO detectors during O3a. Based on the selected segments of data, we then produce a list of valid integer GPS times with the only requirements being that the merger time is at least 100 seconds from the start of the segment (slightly longer than the longest possible waveform), and at least 24 seconds from the end of the segment. These valid GPS times are used for assigning training and validation samples a merger time. The only preprocessing we apply to the data is downsampling it to 2,048 Hz. We compute the PSD for each week of data, which is used for normalizing the SNR time series. We use the Welch method to compute the PSD with a 4-second window for each 1,024-second segment of data. These PSDs are then averaged together for a mean PSD for each week of data.
@@ II.4 Training dataset construction
For our training dataset, we next assign a random GPS time from the first week of O3 to each event. To ensure we have enough unique noise realizations, we use a different random GPS time for each interferometer. We also specify the probability for a sample to contain a glitch. Based on the number of glitches found by Omicron, we chose the Hanford glitch fraction to be 0.1, and the Livingston glitch fraction to be 0.15, as glitches occurred more frequently in Livingston [5]. These glitch fractions are around an order of magnitude higher than the actual rate of glitches during O3, but we specified these fractions to ensure the neural network is sufficiently trained on glitches.
@@
Strain samples with injections are created by first generating the injection waveform with the SpinTaylorT4 approximant [45]. The waveforms are then projected onto the detectors and placed with the merger at the center of a 1024s-long noise segment based on the Hanford GPS time, with the Livingston detector signal offset based on the extrinsic parameters. The strain is then filtered with the 10 selected templates, and the resulting SNR time series are then sliced down to a 2-second time series centered on the merger. As our neural network’s input window is 1 second long, this 2-second slice of SNR time series allows the merger time to be placed randomly within the input window. However, to ensure the entire peak from the merger in both interferometers is completely contained within an SNR time series, we place the merger at least 1/8th of a second from the edges of the window.
@@
Polarization Uniform [0, $\pi$]
Network SNR $\cdots$ $\geq$ 6
@@ III.1 Model architecture
Figure 1: (a) The high-level architecture of our neural network. Each block represents a layer or group of layers, with the height of the block representing the number of convolution kernels and the width representing output size of the layer. The Hanford input branch is shown, and the Livingston input branch has the same architecture. (b) The internal structure of each residual block. (c) A legend of layers and their representation in the diagram. Residual blocks with a purple stripe contain a maximum pooling layer. During training, there is a dropout layer between the dense layers, and a sigmoid activation layer after the last dense layer.
The network is constructed such that it can be split into three subnetworks around the addition layer between the two branches. The three resulting networks (hereafter called the H, L and combiner subnetworks) can then be operated independently. Since the combiner subnetwork is a simple dense network, it can be run much faster on a CPU than the H and L subnetworks that require a GPU to run efficiently. The advantage of this network architecture is that the outputs of the H and L subnetworks can be independently paired up before being input into the combiner subnetwork, which aids in constructing our background to characterize our ranking statistic, as further explained in Sec. IV. One downside of this architecture is that any coincidence information between the two detectors is lost. To mitigate this, an additional input $\Delta t$ is passed to the combiner subnetwork as well as the H and L subnetwork outputs
@@
During training, there are three additional features that are removed during inference. First, we use a dropout layer between each dense layer in the combiner subnetwork to reduce overfitting on the training set. Second, we use a sigmoid activation layer after the final dense layer to constrain the prediction to between 0 and 1. Removing this activation during inference has been shown to be effective at mitigating the resolution limitations of 32-bit precision [53] and allows our ranking statistic to be unbounded. Thirdly, we add a custom layer before the sigmoid activation layer which divides the output of the last dense layer by a factor of 4, which helps prevent the sigmoid layer from rounding predictions to 0 or 1 during training.
@@ III.2 Training
The network was trained using the training and validation datasets described in Sec. II.4. The training and validation samples were weighted in training if their network SNR was greater than 10 by a factor of SNR/10, as we found that the network tended to assign very low prediction values to very high SNR signals, likely because it was mistaking them for glitches. This sample weighting factor also applied to noise-only samples, and so the neural network was additionally penalized for labeling SNR time series with high SNR glitches as signals.
The network was trained with the binary crossentropy loss function using the ADAM optimiser [54], and an initial learning rate of $10^{-4}$. We use the Keras callbacks ReduceLROnPlateau and EarlyStopping to fine-tune the network further: if the validation loss did not improve in 15 epochs, the learning rate was halved, and training stopped if the network’s validation loss did not improve after 25 epochs. We created a custom metric, which we call LogAUC, for EarlyStopping to determine when the network should stop being trained. LogAUC is based on Keras’s AUC metric, which calculates the true alarm probability and false alarm probability at different thresholds (i.e., a receiver operating characteristic, or ROC curve), then computes the area under the resulting curve. Maximizing the area under this curve ensures a trained model is sensitive at the tested false alarm thresholds. LogAUC simply spaces the false alarm rate thresholds logarithmically and calculates the area of the curve in log space, rather than linearly, which ensures that the sensitivity at low false alarm rates is given suitable weight during training. This is important as GW detection requires a focus on low false alarm rates assigned over many orders of magnitude. Training the network typically takes 3 hours on an NVIDIA A100 GPU.
@@ IV Search Method and False Alarm Rate Assignment
Here we describe how we run our search pipeline and compute our ranking statistic, which is used for assigning false alarm rates to events. Since we detect in the SNR time series rather than the strain, our ranking statistic involves several data reduction steps to reduce the computational requirements of the neural network. The ranking statistic for a second of data is computed with the following workflow. First, the SNR time series are computed using the entire template bank. These SNR time series are computed in clusters of 30 due to memory limitations, and are ordered by chirp mass (i.e., all templates in the cluster have a similar chirp mass). For each second of SNR time series, we use a coincident peak-finding algorithm to find the highest network SNR of the time series. A trigger is produced if a peak is found with SNR $>4$ in at least one of the detectors. Only one trigger is saved per SNR time series per second.
@@
The SNR time series with the highest network SNR trigger from each cluster is then used as the input to the neural network. The neural network then makes 16 predictions per second on the SNR time series, i.e., an inference rate of 16 Hz. The moving average of each series of predictions is then used as the ranking statistic on that second of data with that template. Using an inference rate greater than 1 Hz aids with separating glitches and high SNR noise from signal candidates, and this benefit has been noted in previous works [22, 32, 55]. A noise peak is unlikely to produce multiple high-valued predictions in a row, while a peak from a signal would consistently produce high predictions while in the input window. More information on the effect of the inference rate on sensitivity can be found in Sec. V.4. The moving average predictions from each of the template clusters are then compared, and the highest moving average prediction becomes the ranking statistic for that second. In addition to the ranking statistic, the network SNR of the trigger, the time of the peak SNR and the ID of the template that produced the trigger are saved for each second of data.
@@
To characterise our ranking statistic background, we collect the ranking statistic on one week of noise, for which we use the third week of O3. To extend our background, we perform 200 time shifts for each second of data. Time shifts are a simple method for increasing the size of a background, thus making it a more accurate distribution, without requiring an excessive amount of real detector data. Time shifts typically involve shifting the data of two interferometers by a time greater than their light travel time and then computing the ranking statistic of this new pair of data. In our case, however, time shifting the SNR time series would be computationally infeasible, as this would result in a two-hundredfold increase in the number of predictions our neural network would have to make. Our solution to this is to divide the network into the H, L and combiner subnetworks (as described in Sec. III), and time shift the outputs of the H and L subnetworks when collecting our background. The benefit of splitting the model this way is that only the combiner subnetwork has to process 200 time shifts, and the H and L subnetworks only process the single week of background data. Since the combiner subnetwork is much less computationally expensive than the H and L subnetworks, it can predict on 200 time shifted samples when running on a CPU at the same pace as the H and L models running on a GPU. This procedure is similar to that described in [56]. In our implementation, we keep the Hanford SNR time series fixed and shift the Livingston SNR time series by one second for each time shift.
@@
Figure 2: Ranking statistic distribution from 1.8 years of background data, and the corresponding false alarm rates. The orange dashed line shows the Gaussian fit used to assign false alarm rates to ranking statistics higher than the background. The gray region shows the range of the extrapolations assigned using 10 independent subsets of the background.